OpenStudy (anonymous):

A line passes through the point negative four comma one and has a slope of three-halves. If you use this slope to generate another point on the line, what are the coordinates of that point? negative one comma three negative two comma four negative twelve comma two negative eight comma three

OpenStudy (anonymous):

You got to test all the choice out to find the answer for this question

OpenStudy (anonymous):

Slope = gradient and the formula is (y2 - y1) / (x2 - x1)

OpenStudy (anonymous):

So (x1 , y1) is the given coordinates (-4 , 1) and (x2 , y2) are the coordinates from the options

OpenStudy (anonymous):

All you need to do is check every options (y2 - 1) / (x2 - (-4)) = 3 / 2

OpenStudy (anonymous):

so it will equal negative one comma three right

OpenStudy (anonymous):

No

OpenStudy (anonymous):

(3 - 1) / (-1 - (-4)) = 2 / 3 not 3 / 2

OpenStudy (anonymous):

then what shall it be

OpenStudy (anonymous):

Try (-2 , 4) as (x2 , y2)